# Evidence of equilibrium dynamics in human social networks evolving in time

### v2

@@ 1 Introduction
Understanding the dynamics of personal relationships is difficult because these mechanisms operate simultaneously, but also because their behaviours are often coupled, giving rise to effects impossible to predict by analysing them separately. Examples of this include the formation of hierarchies and other combinations such as the ones studied in [30], [22],[31],[32], [33] or [28]. It would thus seem that the combination of this myriad of behaviours would make the evolution of the system unpredictable. However, in this paper we show how, despite all this complexity, the temporal evolution of the network can be described by simple mathematical models and, more importantly, its behaviour can be predicted with very high accuracy. Specifically, we provide evidence that relationship networks exhibit a global behaviour akin to equilibrium dynamics. This means that the macroscopic, average properties of the network – such as the distribution of personal contacts, the prevalence of patterns like edges, triangles, or larger structures, and aggregate metrics like network density – remain constant, fluctuating around stable values, even as individual links continuously evolve at a microscopic level. Interestingly, there are apparent trends of stability in other aspects of human social behaviour, which suggest a natural tendency toward the formation of stable patterns and self-reinforcing dynamics. Thus, Hobbs and Burke show how social connections recover after the death of a friend [34], while Alessandreti et al. found that certain mobility patterns are conserved over time [35], linking these patterns to individuals’ social ties. Other studies, such as [36, 37, 38], stress the stability and robustness of online social behaviours. We will circle back to this research in the discussion section.
@@
Despite its importance, to the best of our knowledge, no prior research has explicitly focused on determining whether social networks are indeed in equilibrium. In this context, we present an example of an empirical social network that is in rigorous equilibrium, given by a rich dataset collected over four years. While this is a single case, we will argue in the discussion that the emergence of equilibrium in social networks can be a very general phenomenon. The importance of this result cannot be overstated. If a social network is in equilibrium, its properties can be predicted more accurately over time, facilitating intervention studies. Interventions can be introduced and compared against the equilibrium network, which serves as a baseline for analysis, helping to strengthen social cohesion or address social isolation. The possibility that this equilibrium behaviour may be general also suggests that certain social behaviours and organization principles transcend cultural, situational or individual differences, connecting individual interactions and broader social structures. From a modelling perspective, equilibrium assumptions enable the use of simpler, more tractable mathematical models, such as statistical mechanics tools, developed for physical systems. Conclusions about the mechanisms driving network dynamics, structural trends, and other insights apply not only to a single observed point in time but to the entire unobserved evolution of the network after the transient formation period. This is specially relevant for methods that employ cross-sectional data to investigate network dynamics, like Exponential Random Graph Models. In fact, these approaches rely on an implicit and often overlooked assumption: the network must be in or near equilibrium, as discussed in [39]. In simple terms, when using cross-sectional data, the researcher assumes that every observation is statistically equivalent to any other, a condition that is fulfilled if the network is in equilibrium.
@@ 2 Results
In this work, we analyse a dataset we have collected between 2020 and 2024 that contains the temporal evolution of the network of personal relationships among 888 people belonging to the Blas de Otero High School in Madrid (see the Materials and Methods section for details on the data collection, data composition and data curation). The reported relationships were coded as +2 (very good), +1 (good), -1 (bad), and -2 (very bad). Almost every student provided this list, resulting in the extraction of a weighted directed network for the entire high school, which we will refer to as snapshot or wave. We repeated this process every 16-20 weeks, and with this data we reconstructed the network at different points in time, collecting 10 snapshots of the network over 4 years. With respect to previous literature, we have introduced some improvements in the data collection process. Respondents are allowed to report an unlimited number of relationships, resulting in richer and more heterogeneous data. We also increased the sample size by one order of magnitude, with a larger number of snapshots, allowing for the detailed study of the dynamics in the long term while maintaining a certain level of granularity. Relationships are self-reported, and they are directed and weighted, between -2 and 2, allowing the consideration of negative relationships. Preliminary analyses of earlier, shorter versions of this dataset have been reported elsewhere [40, 41]. It is worth noting that by coding relationships into these four categories (-2, -1, +1 and +2), we inevitably lose many of the details that characterize social interactions. Thus, the resulting network we analyse in this study is not a direct reflection of the complexity of real personal relationships, but rather a simplified representation. Real-world relationships exist on a spectrum, with layers of emotional, social, and contextual subtleties that cannot be completely captured in these discrete categories. However, even with this simplification, we argue that identifying an equilibrium in this network serves as a meaningful proxy for the underlying dynamics of the real network. The equilibrium we detect in this reduced representation suggests that the macroscopic patterns of stability we observe likely reflect deeper robust social structures. If the simplified relationships follow this equilibrium dynamics, it implies that the more complex relationships they represent also tend toward stability over time.
@@
In total, we surveyed 888 people, but not all of them were present in all waves because the composition of the network changes as time goes by. We surveyed people belonging to a high school over the course of four academic years. Every academic year, new people enter the high school to take the first course, and at the end of the school year, almost all the people belonging to the last course leave the high school (a minor portion of them needs to repeat the course if they fail their subjects). Thus, every academic year, some nodes appear in the network, and others disappear. In Figure 1 we depict this turnover in a visual manner. Interestingly, only 15% of the people are present both in the first and the last waves. This will become a very relevant fact when we show the network is in equilibrium, because this equilibrium will not result from the same group of people interacting over the course of four years, but rather an internal property of the network dynamics independent of the network composition. Due to this turnover, every snapshot of the network contains about 500 active nodes out of the 888.
@@
Drawing an analogy to a social system represented as a network, it is essential to define both the macroscopic properties and the microscopic dynamics. In a social network evolving over time, links appear, disappear, or change in nature. From the perspective of a node, an outgoing or incoming +1 link can become a +2, a -1, a -2 or disappear. We refer to an absent link as a 0 link. This constitutes the microscopic dynamics of the system: the evolution of individual ties/links. From these microscopic components, we can build macroscopic properties of the network. In the network, we define an edge as the connection between two nodes that contains two links, one from the first node to the second and one from the second node to the first. Since we have 5 types of links, -2, -1, 0, +1, +2, there are 25 possible edge types. Notice the distinction between link/tie and edge we introduce in our nomenclature. Although it is usual to treat both as interchangeable, we keep this distinction throughout the paper. Our first macroscopic property is the distribution of different edge types: how many +2+2, +1+2, +1+1, etc., edges exist in the network. Notice that a +2+1 edge is not equivalent to a +1+2, specially from a dynamical perspective. Although both edges can evolve towards a +1-1 edge, in the +2+1 case the +2 needs to become a +1 and the +1 a -1, and in the +1+2 case, the +1 remains unchanged and the +2 becomes a -1. Therefore, we keep this distinction in all computations. A second macroscopic property is the in-degree and out-degree distributions of the nodes. For each type of link (+2, +1, -1, and -2), we count how many nodes have 0 incoming +2 links, 1 incoming +2 links, etc. This process is repeated for outgoing links and every type of link, generating eight degree distributions. A third macroscopic property can be the distribution of different triangle types that can form among three nodes. Other macroscopic properties include structural characteristics of the network, such as clustering, average shortest path length, assortativity, etc.
@@
In this paper, we focus on the assessment of the degree, edge, and triangle distributions since we want to be able to define transition matrices to predict the expected equilibrium states of these macroscopic properties from the dynamics observed, granting the conclusions apply to all macroscopic properties. Before going into the details, it is important to comment on the limitations of our approach from the statistical mechanics point of view. The main challenge is the absence of a continuous concept of time in our analysis. By relying on snapshots taken every 20 weeks, we lose information about the transitions that occur between these intervals. Hence, there is an implicit assumption that no multiple transitions between edge states have occurred within the period between snapshots, minimizing the risk of hidden dynamics. It is clear that this assumption may not hold in all cases: In systems like a gas, for example, relevant changes happen on much shorter timescales, making such large gaps between observations inappropriate. However, in the context of social networks, the underlying dynamics evolve more slowly, and we believe that 20-week intervals provide a sufficient resolution. On the other hand, the use of statistical mechanical techniques in this study necessarily involves certain approximations to simplify the analysis of complex social systems. We aim at striking a balance between revealing meaningful sociological patterns and maintaining the rigour without becoming overly entangled in these mathematical details.
@@
Let us introduce the concept of transition matrix using the edges as our macroscopic property. We define the edges transition matrix $P_{\beta\alpha}$ from snapshot $\alpha$ to snapshot $\beta$ as the matrix whose $ij$ element represents $P(j|i)$, i. e., the conditional probability of an edge of going to the $j$ state in snapshot $\beta$ provided it started in the $i$ state in snapshot $\alpha$ (see Materials and Methods for the construction of this matrix). Additionally, we define the edge state distribution $\pi_{\alpha}$ as a column vector in which each element $i$ is the density of edges of type $i$ in the snapshot $\alpha$. With these two objects, it is direct to see that:
$\pi_{\beta}=P_{\beta\alpha}^{T}\pi_{\alpha}$ (1)
@@
This strategy can also be applied to other macroscopic properties. For instance, at the level of triangles – similar to edges – we can identify all unique triangle states and construct a transition matrix for these states. Given that there are five different types of directed links, the number of combinations grows exponentially with the size of the macroscopic structures analysed. In the case of triangles, we identify 1924 unique triangles (the number we observe, not the theoretical maximum). Thus, we can construct a 1924x1924 transition matrix for triangles and perform a similar test.
@@
Similarly, we can apply this method to degrees. For each node, we have eight types of degrees (in and out degrees for +2, +1, -1, and -2 links). For example, for the +2 in-degree, we examine how many nodes with an initial +2 in-degree transition to a different +2 in-degree in the next wave. This allows us to construct a transition matrix for +2 in-degrees. Although the matrix size varies depending on the specific degree analysed, the strategy remains consistent. We stress that, in all cases, we find that the transition matrices governing the evolution of these macroscopic properties are stationary, indicating that the probabilities driving the network’s evolution are constant over time.
@@
Starting from equation 1, focusing on edges, we can determine whether the system has a stationary state under the dynamics governed by the average transition matrix calculated. If it exists, we can calculate and compare it with the actual state of the network. In equation 1, we obtain the edge abundances in the next wave by multiplying the transpose of the transition matrix by the edge abundances in the current wave. When the system reaches the stationary state, these edge abundances become invariant under the transpose of the transition matrix. In other words, to find the stationary state, we need to solve the following equation:
$\tilde{\pi}=P_{\beta\alpha}^{T}\tilde{\pi}$ (2)
@@
$\tilde{\pi}_{i}(P_{\beta\alpha})_{ij}=\tilde{\pi}_{j}(P_{\beta\alpha})_{ji}$ (4)
Since both $\tilde{\pi}$ and $P_{\beta\alpha}$ are sampled from the data, there is some uncertainty associated with both sides of equation 4. Therefore, the Detailed Balance Condition needs to be fulfilled within the confidence intervals associated with both sides of the equation. Considering this, we find that equation 4 is satisfied for every transition between edge states. There is a small violation of the Detailed Balance Condition in five out of the 325 possible transitions (specifically, going from $+0+0$ to $+2+2$, $+1+2$, $+2+1$, $+1+1$, $+0+1$), but this violation is closely within the confidence intervals. Looking to the magnitude of the violation, we observe that it affects to only the 1.9%, 0.6% 0.4% 1.7% 3.6% of the transitions count involving these links, respectively. Thus, we consider that this violation barely disrupts the equilibrium of the network. It is possible to detect in Figure 3 the effect this violation has in the equilibrium abundances. For the edges and triangles affected, the equilibrium curves are slightly above the empirical abundances, although there is still an almost perfect overlap between both. The method’s ability to detect such a small violation supports the robustness of the equilibrium conditions found for the remaining transitions.

### v4

@@ 1 Introduction
Understanding the dynamics of personal relationships is difficult because these mechanisms operate simultaneously, but also because their behaviours are often coupled, giving rise to effects impossible to predict by analysing them separately. Examples of this include the formation of hierarchies and other combinations such as the ones studied in [30], [22],[31],[32], [33] or [28]. It would thus seem that the combination of this myriad of behaviours would make the evolution of the system unpredictable. However, in this paper we show how, despite all this complexity, the temporal evolution of the network can be described by simple mathematical models and, more importantly, its behaviour can be predicted with very high accuracy. Specifically, we provide evidence that relationship networks exhibit a global behaviour akin to equilibrium dynamics. This means that the macroscopic, average properties of the network – such as the distribution of personal contacts, the prevalence of patterns like edges, triangles, or larger structures, and aggregate metrics like network density – remain constant, fluctuating around stable values, even as individual links continuously evolve at a microscopic level. Interestingly, there are apparent trends of stability in other aspects of human social behaviour, which suggest a natural tendency toward the formation of stable patterns and self-reinforcing dynamics. Thus, Hobbs and Burke show how social connections recover after the death of a friend [34], while Alessandreti et al. found that certain mobility patterns are conserved over time [35], linking these patterns to individuals’ social ties. Other studies, such as [36, 37, 38, 39] stress the stability and robustness of online social behaviours. We will circle back to this research in the discussion section.
@@
Despite its importance, to the best of our knowledge, no prior research has explicitly focused on determining whether social networks are indeed in equilibrium. Our work fits within the broader context of complex systems methods applied to social dynamics, specially from the lens of statistical mechanics [40, 41, 42]. In this context, we present an example of an empirical social network that is in rigorous equilibrium, given by a rich dataset collected over four years. While this is a single case, we will argue in the discussion that the emergence of equilibrium in social networks can be a very general phenomenon. The importance of this result cannot be overstated. If a social network is in equilibrium, its properties can be predicted more accurately over time, facilitating intervention studies. Interventions can be introduced and compared against the equilibrium network, which serves as a baseline for analysis, helping to strengthen social cohesion or address social isolation. The possibility that this equilibrium behaviour may be general also suggests that certain social behaviours and organization principles transcend cultural, situational or individual differences, connecting individual interactions and broader social structures. From a modelling perspective, equilibrium assumptions enable the use of simpler, more tractable mathematical models, such as statistical mechanics tools, developed for physical systems. Conclusions about the mechanisms driving network dynamics, structural trends, and other insights apply not only to a single observed point in time but to the entire unobserved evolution of the network after the transient formation period. This is specially relevant for methods that employ cross-sectional data to investigate network dynamics, like Exponential Random Graph Models. In fact, these approaches rely on an implicit and often overlooked assumption: the network must be in or near equilibrium, as discussed in [43]. In simple terms, when using cross-sectional data, the researcher assumes that every observation is statistically equivalent to any other, a condition that is fulfilled if the network is in equilibrium.
@@ 2 Results
In this work, we analyse a dataset we have collected between 2020 and 2024 that contains the temporal evolution of the network of personal relationships among 888 people belonging to the Blas de Otero High School in Madrid (see the Materials and Methods section for details on the data collection, data composition and data curation). The reported relationships were coded as +2 (very good), +1 (good), -1 (bad), and -2 (very bad). Almost every student provided this list, resulting in the extraction of a weighted directed network for the entire high school, which we will refer to as snapshot or wave. We repeated this process every 16-20 weeks, and with this data we reconstructed the network at different points in time, collecting 10 snapshots of the network over 4 years. With respect to previous literature, we have introduced some improvements in the data collection process. Respondents are allowed to report an unlimited number of relationships, resulting in richer and more heterogeneous data. We also increased the sample size by one order of magnitude, with a larger number of snapshots, allowing for the detailed study of the dynamics in the long term while maintaining a certain level of granularity. Relationships are self-reported, and they are directed and weighted, between -2 and 2, allowing the consideration of negative relationships. Preliminary analyses of earlier, shorter versions of this dataset have been reported elsewhere [44, 45]. It is worth noting that by coding relationships into these four categories ($-2$, $-1$, $+1$, and $+2$), we inevitably lose many of the details that characterize social interactions. Real-world relationships exist on a spectrum, with layers of emotional, social, and contextual subtleties that cannot be completely captured in these discrete categories. Moreover, respondents may not report their relationships with the exact granularity imposed by our discretization, as these relationships are inherently multifactorial and span multiple dimensions of variability. Thus, the resulting network we analyse in this study is not a direct reflection of the complexity of real personal relationships, but rather a simplified representation. Nevertheless, even with this simplification, we argue that identifying an equilibrium in the network serves as a meaningful proxy for the underlying dynamics of real social interactions. The equilibrium we detect in this reduced representation suggests that the macroscopic patterns of stability likely reflect deeper, robust social structures. Importantly, this simplification enables the mathematical rigour of our analysis; if equilibrium behaviour is clearly observable in the simplified model, it implies a genuine stability in the underlying network of relationships, even if the full complexity of these relationships is not entirely captured.
@@
In total, we surveyed 888 people, but not all of them were present in all waves because the composition of the network changes as time goes by. We surveyed people belonging to a high school over the course of four academic years. Every academic year, new people enter the high school to take the first course, and at the end of the school year, almost all the people belonging to the last course leave the high school (a minor portion of them needs to repeat the course if they fail their subjects). Thus, every academic year, some nodes appear in the network, and others disappear. In Figure 1 we depict this turnover in a visual manner. Interestingly, only 15% of the people are present both in the first and the last waves. This will become a very relevant fact when we show the network is in equilibrium, because this equilibrium will not result from the same group of people interacting over the course of four years, but rather an internal property of the network dynamics independent of the network composition. Due to this turnover, every snapshot of the network contains about 500 active nodes out of the 888. Let us define some notation for the discussions that follow. Throughout the paper, we will distinguish between a ‘link’ and an ‘edge’. Although these two concepts are often treated interchangeably, here we differentiate them. A link (or tie) refers to the directional connection from one node to another, whereas an edge represents the pair of links connecting two nodes (one in each direction). We provide a more detailed explanation of this distinction later in the manuscript.
@@
Drawing an analogy to a social system represented as a network, it is essential to define both the macroscopic properties and the microscopic dynamics. In a social network evolving over time, links appear, disappear, or change in nature. From the perspective of a node, an outgoing or incoming +1 link can become a +2, a -1, a -2 or disappear. We refer to an absent link as a 0 link. This constitutes the microscopic dynamics of the system: the evolution of individual ties/links. From these microscopic components, we can build macroscopic properties of the network. In the network, we define an edge as the connection between two nodes that contains two links, one from the first node to the second and one from the second node to the first. Since we have 5 types of links, -2, -1, 0, +1, +2, there are 25 possible edge types. Notice the distinction between link/tie and edge we introduce in our nomenclature. Although it is usual to treat both as interchangeable, we keep this distinction throughout the paper. Our first macroscopic property is the distribution of different edge types: how many +2+2, +1+2, +1+1, etc., edges exist in the network. Notice that a +2+1 edge is not equivalent to a +1+2, specially from a dynamical perspective. Although both edges can evolve towards a +1-1 edge, in the +2+1 case the +2 needs to become a +1 and the +1 a -1, and in the +1+2 case, the +1 remains unchanged and the +2 becomes a -1. Therefore, we keep this distinction in all computations. A second macroscopic property is the in-degree and out-degree distributions of the nodes. For each type of link (+2, +1, -1, and -2), we count how many nodes have 0 incoming +2 links, 1 incoming +2 links, etc. This process is repeated for outgoing links and every type of link, generating eight degree distributions. A third macroscopic property can be the distribution of different triangle types that can form among three nodes. Other macroscopic properties include structural characteristics of the network, such as clustering, average shortest path length, assortativity, centrality metrics, etc.
@@
In this paper, we focus on the assessment of the degree, edge, and triangle distributions since we want to be able to define transition matrices to predict the expected equilibrium states of these macroscopic properties from the dynamics observed, granting the conclusions apply to all macroscopic properties. To assess whether our conclusions extend to more complex structures within the network, we have included an exemplary analysis of betweenness centrality and the details of this supplementary analysis are provided in the Supplementary Material. Before going into the details, it is important to comment on the limitations of our approach from the statistical mechanics point of view. The main challenge is the absence of a continuous concept of time in our analysis. By relying on snapshots taken every 20 weeks, we lose information about the transitions that occur between these intervals. Hence, there is an implicit assumption that no multiple transitions between edge states have occurred within the period between snapshots, minimizing the risk of hidden dynamics. It is clear that this assumption may not hold in all cases: in systems like a gas, for example, relevant changes happen on much shorter timescales, making such large gaps between observations inappropriate. This point is important because there is a close relationship between the timescale of the network dynamics, the measurement intervals, and our ability to detect equilibrium behaviour. In networks with slower dynamics, transient fluctuations may persist longer and result in more frequent apparent violations of the equilibrium, as the system might not have enough time to relax back to equilibrium between observations. Conversely, if the network dynamics are too fast, although the equilibrium state may be accurately detected, the transition matrices might not fully capture the system’s internal dynamics because multiple transitions on the same link could occur between snapshots, distorting the measured change probabilities. However, in the context of social networks, the underlying dynamics evolve more slowly, and we believe that 20-week intervals provide a sufficient resolution. On the other hand, the use of statistical mechanical techniques in this study necessarily involves certain approximations to simplify the analysis of complex social systems. We aim at striking a balance between revealing meaningful sociological patterns and maintaining the rigour without becoming overly entangled in these mathematical details.
@@
Let us introduce the concept of transition matrix using the edges as our macroscopic property. We define the edges transition matrix $m^{K}$ as the matrix whose $ij$ element represents $P(j|i)$, i. e., the conditional probability of an edge of going to the $j$ state in one snapshot provided it started in the $i$ state in the previous snapshot (see Materials and Methods for the construction of this matrix). The letter $K$ labels the transition, such that $m^{A}$ is the transition between snapshots 0 and 1, $m^{B}$ is the transition between snapshots 1 and 2, etc. Additionally, we define the edge state distribution $\pi_{t}$ as a column vector in which each element $i$ is the density of edges of type $i$ in the snapshot $t$. With these two objects, it is direct to see that:
$\pi_{t+1}=(m^{K})^{T}\pi_{t}$ (1)
@@
This strategy can also be applied to other macroscopic properties. For instance, at the level of triangles – similar to edges – we can identify all unique triangle states and construct a transition matrix for these states. Given that there are five different types of directed links, the number of combinations grows exponentially with the size of the macroscopic structures analysed. In the case of triangles, we identify 1924 unique triangles (the number we observe, not the theoretical maximum). Thus, we can construct a 1924x1924 transition matrix for triangles and perform a similar test. Similarly, we can apply this method to degrees. For each node, we have eight types of degrees (in and out degrees for +2, +1, -1, and -2 links). For example, for the +2 in-degree, we examine how many nodes with an initial +2 in-degree transition to a different +2 in-degree in the next wave. This allows us to construct a transition matrix for +2 in-degrees. Although the matrix size varies depending on the specific degree analysed, the strategy remains consistent. We stress that, in all cases, we find that the transition matrices governing the evolution of these macroscopic properties are stationary, indicating that the probabilities driving the network’s evolution are constant over time. In all three cases (edges, triangles, and degrees), we provide the abundances, transition matrices, z-scores, and p-values in the project’s repository as supplementary material.
Figure 4: Comparison between the empirical distributions of the three types of the betweenness centrality considered and theoretical equilibrium distributions computed using the transition matrices. In panel $a)$ it is depicted this comparison for the centrality measured considering only +2 links. In panel $b)$ it is depicted this comparison for the centrality measured considering +2 and +1 links, treating them as equivalent. In panel $c)$ it is depicted this comparison for the centrality measured considering only -1 links. The error-bars correspond to the 95% confidence intervals.
@@
Starting from equation 1, focusing on edges, we can determine whether the system has a stationary state under the dynamics governed by transition matrix calculated. If it exists, we can calculate and compare it with the actual state of the network. In equation 1, we obtain the edge abundances in the next wave by multiplying the transpose of the transition matrix by the edge abundances in the current wave. When the system reaches the stationary state, these edge abundances become invariant under the transpose of the transition matrix. In other words, to find the stationary state associated to a transition matrix $m$ governing the dynamics, we need to solve the following equation:
$\tilde{\pi}=m^{T}\tilde{\pi}$ (2)
@@
$(\pi_{t})_{i}(m^{K})_{ij}=(\pi_{t})_{j}(m^{K})_{ji}$ (4)
Since both $\pi_{t}$ and $m^{K}$ are sampled from the data, there is some uncertainty associated with both sides of equation 4. Therefore, the Detailed Balance Condition needs to be fulfilled within the confidence intervals associated with both sides of the equation (see Materials and Methods for details and the project’s repository for the raw numerical results on these values and confidence intervals). We find that equation 4 is satisfied for almost every transition between edge states within the confidence intervals. We observe small violations of the Detailed Balance Condition in some transitions between edge states. A table summarizing the characteristics of those transitions exhibiting a violation is provided in the Supplementary Material. Overall, the proportion of transitions that do not fulfil the condition is 0.003, 0.01, 0.04, 0.01, 0.023, 0.027, 0.004, 0.01, and 0.007 for transitions 1–9, respectively. In all cases, the magnitude of the violation is minor. Specifically, for the instances where a violation occurs, we have computed the z-scores, defined as the number of standard errors by which the condition is not met. In the majority of cases, the z-score is between 2 and 3, indicating that although the condition is not strictly fulfilled, it is very nearly so. Overall, given the low frequency and small magnitude of these violations, we conclude that they have a negligible effect on the equilibrium of the network. It is possible to detect in Figure 3 the effect this violation has in the equilibrium abundances. For the edges and triangles affected, the equilibrium curves are slightly above the empirical abundances, although there is still an almost perfect overlap between both. The method’s ability to detect such a small violation supports the robustness of the equilibrium conditions found for the remaining transitions. The raw results for the assessment of the detailed balance condition for degrees a triangles are included in the project’s repository as well.
These observed violations might be attributed to simple statistical fluctuations in the network dynamics. Nonetheless, there appears to be a bias in the probability flows involved in these discrepancies favouring the formation of new positive relationships. Although these fluctuations remain compatible with the equilibrium framework we present, it is worthwhile to explore the potential origins of these discrepancies. One plausible explanation is that these deviations may be linked to the fact that we analyse students during a critical developmental stage. During this period, individuals are likely enhancing their cognitive capacities and their ability to manage a larger number of social connections. Being this indeed the case, the equilibrium dynamics framework may be applicable only to short- and medium-term dynamics during these life stages. However, it is important to stress that this interpretation is speculative. The overall results align very closely with the network being in equilibrium, suggesting that, regardless of the underlying cause, these discrepancies have only a minimal effect on the network dynamics.
As a final complementary analysis, we examined whether a metric capturing more complex structural behaviour—specifically, Betweenness Centrality—exhibits a pattern similar to that observed for degrees, edges, and triangles. Remarkably, as we can see in Figure 4, we observe that, in this case as well, the transition matrices are stationary, the detailed balance condition is fulfilled, and the empirical abundances coincide with the theoretical equilibrium distributions (see the Supplementary Material for all technical details). All raw numerical results are available in the project’s repository.